OpenStudy (anonymous):

Write the equation of a line which contains the point (4,3) and (2,5).

OpenStudy (anonymous):

do you know how to find the slope? or do you just need the answer?

OpenStudy (anonymous):

y - 3 = -1(x-4)

OpenStudy (anonymous):

slope formular

OpenStudy (anonymous):

formula*

OpenStudy (anonymous):

do you know how to compute it?

OpenStudy (anonymous):

y = -x +7